# PerturBench: Benchmarking Machine Learning Models for Cellular Perturbation Analysis

### v3

@@ Abstract
We introduce a comprehensive framework for perturbation response modeling in single cells, aimed at standardizing benchmarking in this rapidly evolving field. Our approach includes a modular and user-friendly model development and evaluation platform, a collection of diverse perturbational datasets, and a set of metrics designed to fairly compare models and dissect their performance nuances. Through extensive evaluation of both published and baseline models across diverse datasets, we highlight the limitations of widely used models, such as mode collapse. We also demonstrate the importance of rank metrics which complement traditional model fit measures, such as RMSE, for validating model effectiveness. Notably, our results show that while no single model architecture clearly outperforms others, simpler architectures are generally competitive and scale well with larger datasets. Overall, this benchmarking exercise sets new standards for model evaluation, supports robust model development, and advances the potential of these models to use high-throughput genetic and chemical screens for disease target discovery.
@@ 1 Introduction
Perturbing biological systems such as cells using small molecules and genetic modifications can help researchers to uncover causal drivers of diseases and identify potential therapeutic targets [8, 27, 11]. Advances in CRISPR technology and lab automation have enabled these experiments, which we refer to as perturbation screens, to be conducted at scale with up to hundreds of thousands of perturbations applied in parallel in a single experiment [49]. These perturbation screens have been combined with modern RNA-sequencing technology to measure gene expression profiles at single cell resolution, creating an atlas of cellular snapshots that reveal perturbation effects [36, 54, 37, 1, 15, 8, 51, 61].
@@
However, measuring the perturbation effects of all roughly $20,000$ protein coding genes or $10^{60}$ drug-like chemicals remains prohibitively expensive, especially when combinations of perturbations and multiple tissues/cell types are considered [44, 45]. As a result, there has been a growing interest in generative machine learning approaches that can predict the effects of perturbations on gene expression.
@@
Specifically, researchers have developed models that can generate counterfactual, out of sample (oos) predictions of perturbation effects [19]. One use case, which we call covariate transfer, involves training a model on perturbation effects measured in a set of covariates (e.g., cell lines) and predicting those effects in another covariate where they have not been observed. Another use case — combo prediction, involves training a model on individual perturbation effects and predicting the effects of multiple perturbations in combination. The ultimate goal of these models is to enable in-silico screening across the vast space of unobserved perturbations for novel therapeutic discovery.
Comparing the performance of published models has unfortunately remained a challenge due to inconsistent benchmarking datasets and metrics. The NeurIPS 2023 perturbation prediction challenge provided a novel chemical perturbation, but was limited to just that dataset with the covariate transfer task and a single metric [10]. Additionally, the challenge did not benchmark published models. The sc-perturb database provides datasets with unified metadata, but does not attempt to benchmark models [41]. Ahlmann-Eltze et al., [2], Wenteler et al., [56], Csendes et al., [12], and Wong et al., [58] assessed only the performance of single cell foundation models and GEARS and all studies used only MSE as an evaluation metric, which may not capture key aspects of model performance. Kernfeld et al., [28] focused on using gene regulatory networks to predict perturbation response prediction. Li et al. [31, 32] performed a more comprehensive assessment of published deep learning models across multiple datasets, but were missing strong deep learning baselines and used a limited set of metrics. To our knowledge, none of the existing benchmarking studies offer a model development framework to build perturbation response prediction models. The existing studies also do not perform ablation studies to identify the specific architectural components that drive model performance.
@@
In this work, we provide a rigorous, quantitative assessment of the field of perturbation response modeling by: (1) introducing a highly modular and user-friendly framework in the form of a codebase for model development and evaluation, (2) curating diverse perturbational datasets and defining biologically relevant tasks, (3) defining a set of metrics that enable comparison of different models and capture key failure modes, (4) performing extensive evaluation of published perturbation models, strong baselines, and individual model components. Figure 1 illustrates our approach.
@@
We reproduce key components of published models that cover a spectrum of architectures, and evaluate them against each other and strong baseline models. We specifically test the models on difficult tasks, simulating how they will be deployed in real-life contexts. Our findings reveal that some widely used models are prone to “mode” or “posterior” collapse (see Appendix 9.3 for more details). Since a common use-case of these models is to run in-silico screens that rank perturbations by a desired effect (e.g. reversing a disease state) [55], we propose rank metrics complementary to traditional measures of model fit (e.g. root mean squared error (RMSE)) that specifically assess the models’ ability to accurately order perturbations and act as a safeguards against model collapse. In addition, we demonstrate that models with simple architectures can outperform some of the more sophisticated models, and scale well with larger datasets.
@@ 2 Datasets and Tasks
Many published models have been evaluated on relatively small datasets, where most of the data are seen in the train split. However, in real-world settings, we often have complex datasets that only contain a small fraction of the perturbation effects we are interested in predicting. Thus, we select datasets and create tasks that mirror these real-world challenges. We select five published datasets, Norman et al., [39], Srivatsan et al., [51], Frangieh et al., [18], McFaline-Figueroa et al., [38], and Jiang et al., 2024a [24] which include at least 100 perturbations and cover a diversity of perturbation modalities (chemical vs genetic), combinatorial perturbations, dataset sizes, and covariates. We provide a cursory overview in Table 1 and more information in Appendix 10.1. Here, we define a biological state as a unique set of covariates that we plan to model (e.g. cell line). Details of the preprocessing of the datasets are in the Appendix 10.3.
@@
Dataset Single Dual Modality Biological Cells Task
perturbations perturbations states
@@
We create a covariate transfer split for the Srivatsan20, Frangieh21 and Jiang24 datasets as well as a combo prediction task for the Norman19 dataset. In addition, we study two scenarios: data scaling and imbalanced data. In the former, we benchmark model performance with increasing training data. In the latter, we simulate increasing the imbalance of perturbations observed in different covariates. Details of the experiments are in sections 5.3 and 9.2 respectively. The aim of both scenarios is to simulate how models will be deployed in practice, where there are often complex covariates, imbalanced datasets, and/or large amounts of missing data [see e.g. 16]. Additional details about data splitting implementation can be found in the Appendix 10.4.
@@ 3.1 Modeling counterfactuals
Matching treated outcomes to controls is a common approach to identify treatment effects [see e.g. 52, Section 1.3 for a historical summary]. In the context of perturbation effect prediction, matching control and perturbed cells has been used by a variety of published models such as GEARS [47], scGPT [13], and scFoundation [21]. For the matching approach, ensuring that the control cell is from the same cell type, experiment or batch as the perturbed cells helps reduce potential confounding effects, but cannot account for unobserved sources of variance. A more complex approach is to use optimal transport to identify the control cell most likely to transition into a given perturbed cell [25, 9, 29]. This enables prediction of the full distribution of cellular responses, instead of just the average response.
@@
An alternative to matching methods involves disentanglement [5], which enables models to separate the unperturbed cellular state and the perturbation effect. The compositional perturbation autoencoder (CPA) [35] uses an adversarial classifier to ensure that the unperturbed “basal” state is free of any perturbational information, forcing the perturbation encoder to learn a meaningful representation of the perturbation. These representations are added to control cell encodings during inference to generate counterfactual predictions. Biolord [42] partitions the latent space into subspaces and optimizes those latent spaces to represent covariates and perturbations, which can be recombined during inference to generate counterfactual predictions. sVAE [34] leverages recent results by Lachapelle et al., [30] demonstrating that enforcing a sparsity constraint can induce disentanglement. Bereket and Karaletsos, [6] build on sVAE using an additive conditioning for the perturbations and demonstrate that their model, SAMS-VAE, offers biological interpretability of the latent encodings.
@@ 3.2 Models for benchmarking
In this paper, we implement a range of perturbation response models with diverse architectures such as CPA, Biolord, SAMS-VAE and GEARS. Our aim is to assess the core modeling assumption behind these models, such as the adversarial classifier in CPA and the sparse additive mechanism in SAMS-VAE. See Appendix 10.5 for details. These models are marked with following the model name (e.g. CPA).
To investigate the effect of disentanglement, we ablate the adversarial component from CPA and refer to the new version as CPA (noAdv). The rest of the model is unchanged, that is despite the basal state of cells (i.e. $z_{basal}$) are contaminated with perturbation and covariate information, counterfactual gene expressions are still generated by adding the target perturbation embedding to $z_{basal}$.
@@
We also experiment with scGPT, a single-cell gene expression foundation model [13], to embed gene expression — as inputs to CPA and our latent additive baseline model. Our aim is to understand whether using embeddings from a pretrained foundation model would improve performance.
@@ 3.3 Baseline models
The linear baseline model uses the control matching approach. Given a perturbed cell, $x\textquoteright$ ’, we sample a random control cell with matched covariates, $x$, and reconstruct $x^{\prime}$ by applying one linear layer given the perturbation and covariates:
@@
where $p_{\mathrm{one\_hot}}$ denotes the one-hot encoding of the perturbation and $cov_{\mathrm{one\_hot}}$ denotes one-hot encodings of covariates (e.g. cell type).
@@
Subsequently, we reconstruct the expression value by decoding the added latent space representation $x\textquoteright=f_{\mathrm{dec}}(z_{\mathrm{ctrl}}+z_{\mathrm{pert}})$. All functions $f_{\mathrm{dec}},f_{\mathrm{ctrl}},f_{\mathrm{pert}}$ are implemented as multilayer perceptrons (MLPs) with dropout [50] and layer normalization [4].
@@
We introduce another class of baseline which we refer to as decoder only, that does not leverage gene expression and aims to predict the perturbation effect solely from covariates, or perturbation, or a mix of both. Consequently, model prediction can be expressed as $x^{\prime}=f_{\mathrm{dec}}(z)$ for $z\in\{p_{\mathrm{one\_hot}}\}\cup\{{cov}_{\mathrm{one\_hot}}\}\cup\{(p_{\mathrm{one\_hot}},{cov}_{\mathrm{one\_hot}})\}$. This baseline can be used to establish a performance lower-bound when it does not receive any perturbation information, simulating mode or posterior collapse where the model predicts the same effect for every perturbation.
@@ 4 Benchmarking
Existing perturbation response modeling studies use different metrics, making direct comparison between models difficult. Here, we develop a standardized, modular benchmarking suite with a variety of metrics that mimic key downstream applications and capture common model failure modes. See Appendix 7 for implementation details.
@@ 4.1 Population Aggregation
Since we only have conditionally paired populations of control cells and perturbed cells, a model has to take unperturbed (i.e. control) cell states and predicts what their states would look like if they had been perturbed. To evaluate these counterfactual predictions, we thus compute aggregate measures from the control and perturbed cell populations, and compare the predicted aggregates to the ground truth aggregates. We use the population mean and log fold-change (LogFC) in this paper, but also implement log p-value and distributional metrics such as MMD with RBF kernel and energy distance in our codebase.
@@ 4.2 Metrics
We select a range of metrics that evaluate different aspects of the predicted perturbation response accuracy. We use RMSE (as recommended in Ji et al., 23) to compare the average predicted cell states to the observed cell states. We also use cosine similarity to assess the similarity of the predicted versus observed perturbation effects (LogFCs). However, as we show in Appendix 9.3, these “global” fit-based metrics fail to fully capture all aspects of a model’s performance. Hence, we introduce a set of rank-based metrics that can be seen as a measure of specificity of the model to different perturbations.
@@ 4.2.1 Rank Metrics
Since the space of possible genetic or chemical perturbations is massive, a common application of these models is to rank perturbations by their ability to induce some desired cell state, and select only the top perturbations for experimental validation [55].
Unfortunately, there are no metrics developed to capture this application so far, which results in models that generate predictions with high cosine similarity or low RMSE to the observed gene expression, but fail to capture smaller but key changes that uniquely distinguish the effects of one perturbation from others. In particular, mode or posterior collapse where a model always generates the same prediction irrespective of target perturbation may still result in decent cosine similarity or RMSE on particular dataset.
@@
where $p$ is the number of perturbations that are being modelled, $\hat{x}_{i},x_{i}$ are the predicted and observed (average) expression value of perturbation $i$, and $\mathrm{dist}$ is a generic distance. Figure 2 shows two examples of perturbations predictions and the respective computations of rank metrics. The rank metric is always a number between 00 and $1$, where 00 is a perfect score and $0.5$ is the expected score of a random prediction. Each fit-based metric has a corresponding rank metric.

### v4

@@ Abstract
We introduce a comprehensive framework for modeling single cell transcriptomic responses to perturbations, aimed at standardizing benchmarking in this rapidly evolving field. Our approach includes a modular and user-friendly model development and evaluation platform, a collection of diverse perturbational datasets, and a set of metrics designed to fairly compare models and dissect their performance. Through extensive evaluation of both published and baseline models across diverse datasets, we highlight the limitations of widely used models, such as mode collapse. We also demonstrate the importance of rank metrics which complement traditional model fit measures, such as RMSE, for validating model effectiveness. Notably, our results show that while no single model architecture clearly outperforms others, simpler architectures are generally competitive and scale well with larger datasets. Overall, this benchmarking exercise sets new standards for model evaluation, supports robust model development, and furthers the use of these models to simulate genetic and chemical screens for therapeutic discovery.
@@ 1 Introduction
Perturbing biological systems, such as cells, using small molecules and genetic modifications can help researchers to uncover causal drivers of diseases and identify potential therapeutic targets [9, 26, 11]. Advances in CRISPR technology and lab automation have enabled these experiments, which we refer to as perturbation screens, to be conducted at scale with up to hundreds of thousands of perturbations applied in parallel in a single experiment [49]. These perturbation screens have been combined with modern RNA-sequencing technology to measure gene expression profiles at single cell resolution, creating atlases of cellular snapshots that reveal perturbation effects [35, 54, 36, 1, 15, 9, 51, 61].
@@
However, measuring the perturbation effects of all roughly $20,000$ protein coding genes or $10^{60}$ drug-like chemicals remains prohibitively expensive, especially when taking into account combinations of perturbations and different tissues, cell types, and cell lines [43, 44]. As a result, there has been a growing interest in generative machine learning approaches that can predict the effects of perturbations on gene expression.
@@
Specifically, researchers have developed models that can generate counterfactual, out of sample (oos) predictions of perturbation effects [19]. One use case, which we call covariate transfer, involves training a model on perturbation effects measured in a set of covariates (e.g., cell lines) and predicting those effects in another covariate where the perturbation-covariate pairs have not been observed. combo prediction involves training a model on individual perturbation effects and predicting the effects of multiple perturbations in combination. The ultimate goal is to enable in-silico screens across the vast space of unobserved perturbations to accelerate therapeutic discovery.
Related Works
Comparing the performance of published models has been challenging due to inconsistent benchmarks with different datasets and metrics. The sc-perturb database provides datasets with unified metadata, but does not benchmark models [40]. The NeurIPS 2023 perturbation prediction challenge [53] was a major achievement in standardizing benchmarks, providing a novel chemical perturbation dataset with scRNA-seq readouts measured in PBMCs. The challenge used the covariate transfer task, with metrics including mean RMSE, MAE, and cosine similarity of predicted vs ground truth log p-values. The challenge attracted a large number of submissions, many of which were inspired by published models such as chemCPA [53].
Ahlmann-Eltze et al., [3], Wenteler et al., [56], Csendes et al., [12], and Wong et al., [58] evaluate single-cell foundation models (scFM) such as scGPT, scFoundation, scBERT, Geneformer and UCE, in the context of perturbation response modeling. These studies focus on how these general-purpose models can be fine-tuned for this task, using task-specific models such as GEARS, CPA as well as other baselines (e.g., mean prediction, kNN, random forest, linear models) to highlight the limitations of scFMs for this task. Notably, Wong et al., [58] used our baseline models and rank metric to establish the performance of scGPT and GEARS. These works mostly use well-known model fit metrics such as RMSE/MSE and Pearson correlation between averaged predicted and ground truth expression profiles (Pearson Delta, Pearson LogFC). Wenteler et al., [56] also proposed various distributional metrics, including E-distance, which is equivalent to our energy distance based MMD metric.
Kernfeld et al., [27] systematically assessed a wide array of perturbation response prediction models, with a focus on models that use gene regulatory networks as a form of prior knowledge. A central finding of their work was that simple baselines often matched or outperformed more sophisticated models such as GEARS and Geneformer, which confirm the robust performance of simple approaches in this domain. Recent works by Li et al., 2024a [30] and Li et al., 2024b [31] provide more comprehensive benchmarks of a large set of deep learning models across diverse datasets and metrics. Beyond conventional evaluation, their work introduces novel tasks, such as unseen perturbation/covariate transfer [30, 31] and cell state transition prediction [30]. Notably, while scFMs can excel on unseen perturbation prediction, simpler models often show better performance in the unseen covariate prediction [30].
Contributions
@@
In this work, we (1) introduce a highly modular and user-friendly framework in the form of a codebase (Perturbench) for model development and evaluation, (2) curate diverse perturbational datasets and define biologically relevant tasks, (3) develop metrics that enable rigorous model comparison and capture key failure modes, and (4) perform extensive evaluation of published perturbation models, strong baselines, and individual model components. Figure 1 illustrates our approach.
We reproduce key components of published models that cover a spectrum of architectures, and evaluate them alongside strong baseline models. We specifically test the models on difficult tasks, simulating how they will be deployed in real-life contexts. Our findings reveal that some widely used models are prone to “mode” or “posterior” collapse (see Appendix C.7 for more details). Since a common use-case of these models is to run in-silico screens that rank perturbations by a desired effect (e.g. reversing a disease state) [55], we propose rank metrics complementary to traditional measures of model fit (e.g. root mean squared error (RMSE)) that specifically assess the models’ ability to accurately order perturbations and detect model collapse. In addition, we demonstrate that models with simple architectures can outperform more sophisticated models when trained on larger datasets.
@@ 2 Datasets and Tasks
Many published models have been evaluated on relatively small datasets, where most of the data are seen in the train split. However, in real-world settings, we often have complex datasets that only contain a small fraction of the perturbation effects we are interested in predicting. Thus, we select datasets and create tasks that mirror these real-world challenges. We select six published datasets, Norman et al., [38], Srivatsan et al., [51], Frangieh et al., [18], McFaline-Figueroa et al., [37], Jiang et al., 2024a [23], and Szałata et al., [53] (OP3) which include at least 100 perturbations and cover a diversity of perturbation modalities (chemical vs genetic), combinatorial perturbations, dataset sizes, and covariates. We provide a cursory overview in Table 1 and more information in Appendix D.1. Here, we define a biological state as a unique set of covariates that we plan to model (e.g. cell type/line). Dataset preprocessing details can be found in Appendix D.3.
@@
Dataset Single Dual Modality Primary Biological Cells Task
pert. pert. cells states
@@
OP3 144 0 chemical 4 296,147 covariate transfer
We create covariate transfer tasks for the Srivatsan20, Frangieh21, Jiang24 and OP3 datasets as well as a combo prediction task for the Norman19 dataset. In addition, we study two scenarios: data scaling and imbalanced data. In the former, we benchmark model performance with increasing training data. In the latter, we simulate increasing the imbalance of perturbations observed in different covariates. Details of the experiments are in sections 5.3 and C.6 respectively. The aim of both scenarios is to simulate how models will be deployed in practice, where there are often complex covariates, imbalanced datasets, and/or large amounts of missing data [see e.g. 16]. Additional details about data splitting implementation can be found in the Appendix D.4.
@@ 3.1 Modeling counterfactuals
Matching treated outcomes to controls is a common approach to identify treatment effects [see e.g. 52, Section 1.3 for a historical summary]. In the context of perturbation effect prediction, matching control and perturbed cells has been used by a variety of published models such as GEARS [46], scGPT [13], and scFoundation [20]. For the matching approach, ensuring that the control cell is from the same cell type/line, experiment or batch as the perturbed cells helps reduce potential confounding effects, but cannot account for unobserved sources of variance. A more complex approach is to use optimal transport to identify the control cell most likely to transition into a given perturbed cell [24, 10, 28]. This enables prediction of the full distribution of cellular responses, instead of just the average response.
@@
An alternative to matching methods involves disentanglement [6], which enables models to separate the unperturbed cellular state and the perturbation effect. The compositional perturbation autoencoder (CPA) [34] uses an adversarial classifier to ensure that the unperturbed “basal” state is free of any perturbational information, forcing the perturbation encoder to learn a meaningful representation of the perturbation. These representations are added to control cell encodings during inference to generate counterfactual predictions. Biolord [41] partitions the latent space into subspaces and optimizes those latent spaces to represent covariates and perturbations, which can be recombined during inference to generate counterfactual predictions. sVAE [33] leverages recent results by Lachapelle et al., [29] demonstrating that enforcing a sparsity constraint can induce disentanglement. Bereket and Karaletsos, [7] build on sVAE using an additive conditioning for the perturbations, leading to SAMS-VAE which has biologically interpretable latent encodings.
@@ 3.2 Models for benchmarking
In this paper, we implement a range of perturbation response models with diverse architectures such as CPA, Biolord, SAMS-VAE and GEARS. Our aim is to assess the core modeling assumption behind these models, such as the adversarial classifier in CPA and the sparse additive mechanism in SAMS-VAE. See Appendix D.5 for details. These models are marked with following the model name (e.g. CPA).
To investigate the effect of disentanglement, we ablate the adversarial component from CPA and refer to the new version as CPA (noAdv). The rest of the model is unchanged. Despite the basal state of cells (i.e. $z_{basal}$) are contaminated with perturbation and covariate information, counterfactual gene expressions are still generated by adding the target perturbation embedding to $z_{basal}$. We verified that our CPA implementation performs at least as well as the published version in C.7.4.
@@
We also experiment with scGPT, a single-cell gene expression foundation model [13], to embed gene expression — as inputs to CPA and our Latent Additive baseline model. Our aim is to understand whether using cell embeddings from a pretrained foundation model improve performance.
@@ 3.3 Baseline models
The linear baseline model uses the control matching approach. Given a perturbed cell, $x$’, we sample a random control cell with matched covariates, $x$, and reconstruct $x^{\prime}$ by applying one linear layer given the perturbation and covariates:
@@
where $p_{\mathrm{one\_hot}}$ denotes the one-hot encoding of the perturbation and $cov_{\mathrm{one\_hot}}$ denotes one-hot encodings of covariates (e.g. cell type/line).
@@
Subsequently, we reconstruct the expression value by decoding the added latent space representation $x’=f_{\mathrm{dec}}(z_{\mathrm{ctrl}}+z_{\mathrm{pert}})$. All functions $f_{\mathrm{dec}},f_{\mathrm{ctrl}},f_{\mathrm{pert}}$ are implemented as multilayer perceptrons (MLPs) with dropout [50] and layer normalization [5].
@@
We introduce another class of baseline which we refer to as Decoder-Only, that does not leverage gene expression and aims to predict the perturbation effect solely from covariates, or perturbation, or a mix of both. Consequently, model prediction can be expressed as $x^{\prime}=f_{\mathrm{dec}}(z)$ for $z\in\{p_{\mathrm{one\_hot}}\}\cup\{{cov}_{\mathrm{one\_hot}}\}\cup\{(p_{\mathrm{one\_hot}},{cov}_{\mathrm{one\_hot}})\}$. This baseline can be used to establish a performance lower-bound when it does not receive any perturbation information, simulating mode or posterior collapse where the model predicts the same effect for every perturbation.
@@ 4 Benchmarking
Existing perturbation response modeling studies use different metrics, making direct comparison between models difficult. Here, we develop a standardized, modular benchmarking suite with a variety of metrics that mimic key downstream applications and capture common model failure modes. See Appendix A for implementation details.
@@ 4.1 Population Aggregation
Since we only have conditionally paired populations of control cells and perturbed cells, a model has to take unperturbed (i.e. control) cell states and predicts what their states would look like if they had been perturbed. To evaluate these counterfactual predictions, we thus compute aggregate measures from the control and perturbed cell populations, and compare the predicted aggregates to the ground truth aggregates. We use the population mean and log fold-change (LogFC) to assess the average response to perturbation, and compute distributional metrics such as maximum mean discrepancy (MMD) and differential gene expression (DEG) t-scores to assess the full distribution of responses.
@@ 4.2 Metrics
We select a range of metrics that evaluate different aspects of the predicted perturbation response accuracy. We use RMSE (as recommended in Ji et al., 22) to compare the average predicted cell states to the observed cell states and cosine similarity to assess the fidelity of predicted versus observed effects (LogFCs). However, as we show in Appendix C.7, these “global” fit-based metrics fail to fully capture all aspects of a model’s performance. Hence, we introduce a set of rank-based metrics that can be seen as a measure of specificity of the model to different perturbations.
@@ 4.2.1 Rank Metrics
Since the space of possible genetic or chemical perturbations is massive, a common application of these models is to rank perturbations by by their ability to induce some desired cell state, and select only the top perturbations for experimental validation [55].
Unfortunately, there are no metrics developed for perturbation ordering, which results in models that generate predictions with high cosine similarity or low RMSE to the observed gene expression, but fail to capture smaller but key changes that uniquely distinguish the effects of one perturbation from others. In particular, mode or posterior collapse where a model always generates the same prediction irrespective of target perturbation may still result in decent cosine similarity or RMSE on particular dataset.
@@
where $p$ is the number of perturbations that are being modelled, $\hat{x}_{i},x_{i}$ are the predicted and observed (average) expression value of perturbation $i$, and $\mathrm{dist}$ is a generic distance. Figure 2 shows two examples of perturbations predictions and the respective computations of rank metrics. The rank metric is always a number between 0 and $1$, where 0 is a perfect score and $0.5$ is the expected score of a random prediction. Each fit-based metric has a corresponding rank metric. We verified the calibration of this rank metric in Appendix C.7.5.
4.2.2 Distributional Metrics
We evaluate model predictions with metrics that capture the full distribution of perturbation response across cells, such as MMD and differentially expressed gene (DEG) recall.
Our MMD metric computes the energy distance between predicted and the ground truth perturbed cells, either in the gene expression or principal component analysis (PCA) space. To embed cells with PCA, we fit a PCA model to the ground truth test split, retaining the top 256 PCs, and project predicted cells onto the ground truth PCs. DEG recall measures the fraction of ground truth DEGs that can be recalled from the predicted cells. DEGs for the predicted and ground truth perturbed cells are computed relative to the control cells, using the scanpy’s tl.rank_genes_groups method with default parameters  [57]. The t-scores are sorted and the top 20 are retained for comparison. We leave DecoderOnly DEG recall empty as it cannot model variance in perturbation response.